I am so confused. I don't understand how to do the following problem:
4x-2y=12 directions say Plot the graphs of the following given equations.Use the "X" and "Y" intercept method, and a check point. Use the given "X" value when solving for the check point. Really confused please help me understand.

Answer by rfer(16322)

y int is when x=0
x int is when y=0
4x-2y=12
4*0-2y=12
-2y=12
y=-6
(0,-6)
-----------------------
4x-2*0=12
4x=12
x=3
(3,0)
-----------------
you did not give the check point so I picked one x=2
4*2-2y=12
8-2y=12
-2y=4
y=-2
(2,-2)